OpenStudy (anonymous):

what are the zeros of the function f(x)=(x+5)(2x-8)

geerky42 (geerky42):

Set f(x) equal to 0, then solve for x.

geerky42 (geerky42):

\[(x+5)(2x-8)=0\] So either \(x+5=0\) or \(2x-8=0\) You have two equations to solve for x.

OpenStudy (anonymous):

x=-5 x=4 thanks!